OpenStudy (anonymous):

(a) Let F = (4x + 3y) i  + (3y + 8z) j  + (8z + 4x)k Let C be the circumference of the circle of radius 8 in the plane x + y + z = 5, oriented counterclockwise as viewed from the point (5, 5, 5). Find integral_C F · dr = (b) Now suppose we know G = (4x^2 + 3y^2) i  + (3y^2 + 8z^2) j  + (8z^2 + 4x^2)k but only very near to the point (5, 5, 5). Let R be a regular octagon with area 0.08 in the plane y = 5 and C be its perimeter oriented counterclockwise when viewed from (0, 10, 0). Estimate integral_C G · dr≈ Give your answer to 3 decimal places.

OpenStudy (anonymous):

@ganeshie8 do you think the best way to find the integral is to use parametric eqn?

OpenStudy (anonymous):

or can I use stokes' thm?

ganeshie8 (ganeshie8):

you have a closed path so yes you can use stokes

OpenStudy (anonymous):

ok I get a little confuse with stokes thm, do i just find the curl and multiply it by the area of the circle? @ganeshie8

ganeshie8 (ganeshie8):

what does stokes thm say ?

OpenStudy (chosenmatt):

@ganeshie8 is helping you he's really good at math!

OpenStudy (anonymous):

curlF dot dA

OpenStudy (anonymous):

I know he is :)

OpenStudy (anonymous):

curlF= -8, -4, -3

ganeshie8 (ganeshie8):

\[\oint_C \vec{F}\cdot d\vec{r} = \iint _S \mathrm{curl}(\vec{F} )\cdot \hat{n} dS\]

ganeshie8 (ganeshie8):

find ndS and take the dot product with curl

OpenStudy (anonymous):

whats ndS? the normal?

ganeshie8 (ganeshie8):

yes normal to the plane

OpenStudy (anonymous):

so <5,5,5>

ganeshie8 (ganeshie8):

\[\hat{n}dS = \dfrac{\vec{N}}{\vec{N}\cdot \hat{k}}dxdy\]

ganeshie8 (ganeshie8):

\[\hat{n}dS = \dfrac{\langle 5,5,5 \rangle }{\langle 5,5,5 \rangle\cdot \langle 0,0,1 \rangle}dxdy\]

OpenStudy (anonymous):

so I don't need the area of the circle pi(8)^2?

ganeshie8 (ganeshie8):

\[\hat{n}dS = \langle 1,1,1 \rangle dxdy\]

OpenStudy (anonymous):

so its just 5?

ganeshie8 (ganeshie8):

plug that in stokes thm formula

OpenStudy (anonymous):

i'm confuse, how do you ge <1,1,1>

OpenStudy (anonymous):

I'm getting -15 @ganeshie8

OpenStudy (anonymous):

do i multiply -15 times pi(8)^2

ganeshie8 (ganeshie8):

\[\begin{align}\oint_C \vec{F}\cdot d\vec{r} &= \iint _S \mathrm{curl}(\vec{F} )\cdot \hat{n} dS \\~\\ &= \iint_S \langle -8,-4,-3 \rangle \cdot \langle 1,1,1 \rangle dxdy \\~\\ &= \iint_S -15 dxdy \\~\\ &= -15\iint_S 1 dxdy \\~\\ \end{align}\]

OpenStudy (anonymous):

so 1dxdy is the area? @ganeshie8

ganeshie8 (ganeshie8):

yes but it is the area of projection of circular disk onto the xy plane

ganeshie8 (ganeshie8):

\[\begin{align}\oint_C \vec{F}\cdot d\vec{r} &= \iint _S \mathrm{curl}(\vec{F} )\cdot \hat{n} dS \\~\\ &= \iint_S \langle -8,-4,-3 \rangle \cdot \langle 1,1,1 \rangle dxdy \\~\\ &= \iint_S -15 dxdy \\~\\ &= -15\iint_S 1 dxdy \\~\\ &= -15 (\text{Area of projection of circular disk onto xy plane })\\~\\ &= -15(\pi \times 8^2 \times \frac{1}{\sqrt{3}}) \end{align}\]

ganeshie8 (ganeshie8):

you dont have answers for these problems ?

ganeshie8 (ganeshie8):

i would like @eliassaab sir to double my work

ganeshie8 (ganeshie8):

or @Kainui @SithsAndGiggles

OpenStudy (anonymous):

no I don't :/ and I can't try it out because it was due tusday

ganeshie8 (ganeshie8):

its webassign right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

this problem its hard, I have like no idea what to do for part b @ganeshie8

ganeshie8 (ganeshie8):

part b should be easy if you understand part a

ganeshie8 (ganeshie8):

for part b : find the curl of G at (5,5,5) and dot with ndS

OpenStudy (anonymous):

so I find the curl first and than sub in (5,5,5)?

OpenStudy (anonymous):

cool tha's what I got, i just need to multiply it by (0.08) ? @ganeshie8

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

G = (4x^2 +3y^2)i+(3y^2+ 8z^2)j+(8z^2+ 4x^2)k curl(G) = <-16z, -8x, -6y> curl(G) at (5,5,5) = <-80, -40, -30> \[\int_C \vec{G}\cdot d\vec{r} = \iint_S \mathrm{curl}(G) \cdot \hat{n} dS \approx \iint_S \langle -80,-40,-30\rangle \cdot \langle 0,1,0\rangle dxdz\]

OpenStudy (anonymous):

so just just (-40) * (.08) ? @ganeshie8

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

-3.200

OpenStudy (anonymous):

wondering why they would ask for 3 decimal places

ganeshie8 (ganeshie8):

yeah let me know once you have the answers

OpenStudy (anonymous):

i get -3.2

OpenStudy (anonymous):

@ganeshie8 to bad that I can't check the answer :/

OpenStudy (anonymous):

should the answer be negative?

ganeshie8 (ganeshie8):

yes our orientation and normal vector look good

OpenStudy (anonymous):

I guess its time to more to a new question haha thank you so much for your help :) you are the best